OpenStudy (otonogold):

Divide and simplify if possible. square root 250 x^16 over square root 2x

OpenStudy (agent0smith):

Break down 250 = 2*5*5*5 so you can simplify it

OpenStudy (otonogold):

sqrt 250 / sqrt 2 = sqrt 250/2 = sqrt 125 = sqrt 25*5 = sqrt 25 * sqrt 5= 5* sqrt 5= 5 sqrt 5

OpenStudy (agent0smith):

actually do this first\[\Large \frac{ \sqrt{ 250x^{16}} }{ \sqrt{2x} } =\sqrt{\frac{ 250x^{16} }{ 2x } }\]

OpenStudy (otonogold):

sqrt 250 * sqrt 2 and sqrt x^16 / sqrt x ? divide that

OpenStudy (agent0smith):

\[\Large 5 \sqrt5*\sqrt{\frac{ x^{16} }{ x } }\]

OpenStudy (otonogold):

not multiplication I meant to put a / lol

OpenStudy (skullpatrol):

$$\Huge (ab)^m=a^mb^m$$ $$\Huge ({\frac{a}{b}})^m=\frac{a^m}{b^m}$$

OpenStudy (otonogold):

nvm -.-

OpenStudy (agent0smith):

\[\Large 5 \sqrt5*\sqrt{x^{15} } = 5 \sqrt5*\sqrt{x^{14}*x }\]

OpenStudy (otonogold):

we don't know if x>0. If we know that x>0, then we can safely simplify by doing things like x2−−√=x but consider what happens if x<0 ..

OpenStudy (agent0smith):

You have to put abs. value signs.

OpenStudy (agent0smith):

x=0 would also not be a possible value in the original problem, since there's an x in the denominator.

OpenStudy (otonogold):

-.- This is aggravating.. so I have to solve the equation you just put?

OpenStudy (agent0smith):

Just simplify this\[\Large 5 \sqrt5*\sqrt{x^{14}}*\sqrt x \]

OpenStudy (agent0smith):

is that a new question...?

OpenStudy (otonogold):

so typo so I got 5 x^7 sqrt(5x)

OpenStudy (agent0smith):

Yeah, but you should put abs value signs on x^7

OpenStudy (otonogold):

What do you mean?

OpenStudy (agent0smith):

\[\Large 5 \left| x^7 \right| \sqrt{5x}\]